OpenStudy (anonymous):

a bag contains 6 red marbles, 6 white marbles, and 4 blue marbles find P (red or blue)

OpenStudy (anonymous):

the ratio of the total of red and blue marbles to the total number of marbles

OpenStudy (anonymous):

how many marbles total?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

right, and the total of red and blue together?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

ok so \(\frac{12}{16}\) which no doubt reduces

OpenStudy (anonymous):

actually that is wrong, i count 10

OpenStudy (anonymous):

sorry yeah 10

OpenStudy (anonymous):

6 red. 4 blue, total is 10 so \(\frac{10}{16}=\frac{5}{8}\)

OpenStudy (anonymous):

hope it was clear, and also clear that it was not too hard once you know what to do